# Searching for GeV Gamma-Ray Polarization and Axion-Like Particles with AMS-02

### v1

@@ V.2 Galactic sources
Figure 10: Left panel: region of ALP parameter space where the expected confidence level for detection of ALP-induced polarization (assuming initially unpolarized sources) exceeds $95\%$, for ten years of data from AMS-02 (red) and AMS-100 (blue). The sources employed are four Galactic supernova remnants (see text for details). Right panel: region of ALP parameter space where the expected confidence level for detection of ALP-induced polarization exceeds $95\%$, for ten years of data from AMS-100. Different lines correspond to different source populations, + (in the case of pulsars) different assumptions for the initial polarization. The two cases with $100\%$ initial linear polarization correspond respectively to the polarization at the source being aligned with the transverse Galactic magnetic field (green dotted line) and the angle between the two being $\approx\tan^{-1}3$ (purple solid line), which maximizes the initial oscillation amplitude.
We summarize the regions of ALP parameter space where we expect to be able to exclude the relevant null hypotheses, using observations of Galactic sources, in Fig. 10. We show the $95\%$ confidence sensitivity to polarization for AMS-02 and AMS-100 in the left panel, and in the right panel show the sensitivity to energy-dependence of polarization for a range of Galactic sources and assumptions about the initial polarization. In particular, we find that the contours for pulsars and supernova remnants are rather similar, although unpolarized (or circularly polarized) sources allow for probing slightly lower $g_{a\gamma\gamma}$.
@@
Figure 11: Left panel: Region with expected confidence level $95\%$ or higher for detection of non-vanishing linear polarization, compared to other existing constraints. Right panel: Region with expected confidence level $95\%$ or higher for detection of energy dependence in linear polarization, compared to other existing constraints. Again, the possible signal from Fermi-LAT [15] is shaded by light blue. The other existing constraints are plotted using AxionLimitPlotter [55].
@@
Similarly, a null result in this search (even with AMS-100) would not nominally constrain new ALP parameter space due to the overlap with existing bounds, but could help test models which evade these bounds due to specific model-dependent features. The uncertainty in the $B$-field modeling is less severe in this case than for the NGC1275 study, especially as there are multiple sources for which comparable sensitivity would be expected (see App. B), but would need to be accounted for in a detailed study of the implications of either a detection or a null result.
@@ V.4 The possibility of a Galactic supernova
The field of view for AMS-02 conversion photons can be estimated as 0.5-0.7 sr (depending mildly on energy), from the results for acceptance and effective area provided in Ref. [24], so the chance that a given Galactic supernova occurs within the field of view is only $\mathcal{O}(5\%)$ (we caution that this is only an order-of-magnitude estimate that does not account for AMS-02’s non-uniform exposure). Nonetheless, we will here take the optimistic approach of assuming a Galactic supernova at a distance of 10 kpc, located such that its photons reach AMS-02 with near-perpendicular incidence. To estimate the counts from such an event, we employ the example axion spectrum arising from pions (which is the dominant contribution), integrated for 10s after the supernova, calculated by Ref. [56] for $g_{a\gamma\gamma}=10^{-12}\text{GeV}^{-1}$ and $E=0.1-1$ GeV. We weight this spectrum by the effective area for AMS-02 pair-conversion photons given in Ref. [24], for $E=0.2-1$ GeV (the effective area falls rapidly to zero below 0.2 GeV). We find that with the “KSVZ” parameters of Ref. [56], we would expect roughly $9.5\times 10^{10}(g_{a\gamma\gamma}/10^{-12}\text{GeV})^{2}$ axions incident on the detector; the measurable photon counts will be this number multiplied by the axion-photon conversion probability. For the “ALP” parameters of that reference (differing from the“KSVZ” case in the couplings relevant for axion production in the supernova), the corresponding number is roughly $26000(g_{a\gamma\gamma}/10^{-12}\text{GeV})^{2}$ axions.
@@
With the red supergiant conversion probabilities, we see that in order to distinguish a fully polarized signal from an unpolarized signal at 95% confidence (i.e. requiring MDP $<1$, so at least $\sim 600$ counts), for AMS-02 we would need $g_{a\gamma\gamma}\gtrsim 7\times 10^{-12}$ GeV-1, for the ALP benchmark and for masses $m_{a}\lesssim 2\times 10^{-11}$ eV (where the conversion in galactic fields dominates). For larger masses, up to $m_{a}\lesssim 5\times 10^{-5}$ eV, the corresponding sensitivity would be $g_{a\gamma\gamma}\gtrsim 3\times 10^{-11}$ GeV-1. For the KSVZ benchmark, the corresponding numbers are $g_{a\gamma\gamma}\gtrsim 2\times 10^{-13}$ GeV-1 ($m_{a}\lesssim 2\times 10^{-11}$ eV) and $g_{a\gamma\gamma}\gtrsim 7\times 10^{-13}$ GeV-1 ($m_{a}\lesssim 5\times 10^{-5}$ eV).
Since the axion luminosity and the conversion probability each scale independently as $g_{a\gamma\gamma}^{2}$, an AMS-100-like experiment would improve the sensitivity to $g_{a\gamma\gamma}$ by a factor of $100^{1/4}\approx 3$. A larger field of view could also increase the probability of seeing the supernova in the first place. Similarly, if the stellar conversion probability we use above is underestimated by roughly an order of magnitude (as might be the case for a blue supergiant progenitor with kG surface magnetic field), that would improve the sensitivity to $g_{a\gamma\gamma}$ by a factor of $10^{1/4}\approx 1.8$ in the high-ALP-mass region where conversion on stellar magnetic fields dominates.
@@ VI Summary and Outlook
We have analyzed the detectability of gamma-ray polarization, both astrophysical and induced by new physics such as ALP-photon mixing, in the AMS-02 detector and possible successor experiments. Compared with existing X-ray polarization measurements, gamma-ray polarization would both provide a new window on astrophysical production mechanisms for gamma-rays in bright sources, and could offer sensitivity to regions of ALP parameter space where the critical energy for oscillations is high and consequently the signal is suppressed at low photon energies.
We have shown that observations of the bright AGN NGC2175 with AMS-02 cannot set any ($95\%$ confidence) limits on ALP parameter space in the absence of a detection of polarization, due to uncertainties in the $B$-field modeling. Within our framework for $B$-field modeling, for the median $B$-field (ranked by significance of ALP signal) AMS-02 would only have sensitivity (defined as an expected $p$-value for a signal of 0.05 or less) to parameter space that is already nominally excluded by competing constraints; sensitivity to un-excluded parameter space would rely on more optimistic $B$-field models. However, a successor experiment with increased exposure (by a factor of 100) would extend the $g_{a\gamma\gamma}$ reach down to $\sim 8\times 10^{-13}$ GeV-1 for masses up to $\sim 2\times 10^{-9}$ eV, with the median $B$-field configuration, and would test new parameter space even for pessimistic $B$-field configurations (at the $5$th percentile of significance, among $B$-field models we tested). Such an experiment could also detect polarization variation with respect to energy, in a similar parameter region to where Fermi-LAT can search for intensity variation with respect to energy; a direct comparison of the sensitivity requires a combined analysis with a common $B$-field model, and would also depend on the exact properties of such a successor experiment.
@@
Galactic sources appear less favorable for probing small $g_{a\gamma\gamma}$ due to the smaller propagation distance through substantial $B$-fields (compared to NGC1275, which is located in the core of the Perseus cluster). We have shown that the region where AMS-02, or even AMS-100, has sensitivity is nominally excluded by existing bounds from CAST, SHAFT, and ABRA 10cm. This is broadly consistent with previous studies of energy-dependent modulation of the intensity signal in gamma-rays from pulsars, which claimed a nominal detection of a signal, but in ALP parameter space that is in tension with existing bounds [15]. Some models have been proposed to alleviate or remove these tensions, and in such models, the AMS-100 sensitivity region from our work overlaps with the signal region found in the literature for intensity fluctuations, and could thus serve as an independent check for at least part of the parameter space. Similarly, gamma-ray observations of a Galactic supernova could offer sensitivity to new regions of ALP parameter space [56], and we have estimated the sensitivity of AMS-02 to the polarization signal from such a source (assuming the roughly $5\%$ chance that the supernova occurs within the AMS-02 field of view). Polarization is unlikely to be the first detection channel for ALPs in such an event, as a polarization measurement requires a substantial photon flux (in AMS-02-like detectors, at least 600 photons are required for a fully polarized signal to be above the MDP with $p=0.05$), but could serve as a useful cross-check on the ALP origin of such a signal.
@@ Acknowledgements
We thank Francesca Calore and Ben Safdi for helpful conversations, and Carlos Maña Barrera for his work on measuring polarization in AMS-02. YJ is grateful for the support of DOE including resources from the National Energy Research Scientific Computing Center under Contract No. DE-AC02- 05CH11231. XZ is supported by DOE grant number DE-SC0024112. TRS was supported in part by a Guggenheim Fellowship; the Edward, Frances, and Shirley B. Daniels Fellowship of the Harvard Radcliffe Institute; the Bershadsky Distinguished Fellowship of the Harvard Physics Department; and the Simons Foundation (Grant Number 929255, T.R.S). TRS thanks the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Mainz Institute for Theoretical Physics for their hospitality during the completion of this work; this research was supported in part by grant no. NSF PHY-2309135 to KITP, and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This work was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567.
@@ Appendix A Note on choice of energy binning
We demonstrate this point in Fig. 12 and Fig. 13, which shows the sensitivity regions for the energy-dependence analysis for NGC1275 and the brightest galactic sources for the cases of 10 bins and 4 bins. We see that for galactic sources, adding more energy bins does nothing more than slightly filling the “holes” in sensitivity. For NGC1275, when the median-$p$-value $B$-field configuration is chosen for each parameter point (underlying contours), the binning makes very little difference. But when we instead choose a “pessimistic” $B$-field configuration (ranked 95th out of 100 in terms of $p$-value at that point), using only 4 bins almost completely eliminates the region where the expected confidence level for detection exceeds $95\%$, in particular erasing sensitivity at lower $m_{a}$, since the “holes” in sensitivity are exactly what dominates the “pessimistic” case; thus a finer binning improves sensitivity quite considerably. This justifies our choice to use only 4 bins for the Galactic sources for the sake of computing resources, but 10 bins for NGC1275 to compensate for the large uncertainties coming from our ignorance of the B-field configuration. In a real analysis, where there is only one dataset (rather than a large ensemble of simulations), a finer binning (limited by the AMS-02 energy resolution) might be appropriate.
@@
Figure 12: Expected confidence level for detection (exclusion of the null hypothesis) with ten years data of AMS-100 from NGC1275. Contours show the results for the median $B$-field configuration (ranked by significance); the cross-hatched region corresponds to greater than $95\%$ confidence level ($p\leq 0.05$) for a pessimistic $B$-field model (95th/100 in significance). Left panel: results of analysis with 4 log-spaced energy bins. Right panel: results of analysis with 10 log-spaced energy bins, as in Fig. 8.
@@
Figure 13: Sensitivity to energy-dependence of polarization with 10 years of AMS-100 data, as in the right panel of Fig. 10. Left panel: contours of expected confidence level $95\%$ ($p=0.05$), for several individual Galactic sources with 4 log-spaced energy bins. Right panel: results of analysis with 10 log-spaced energy bins.
Appendix B Impact of including or omitting Galactic sources
For Galactic sources, combining polarization detections from multiple sources could increase the significance of any signal and serve as a non-trivial test for the ALP hypothesis. While for any one source the Earth may accidentally lie at a point where the ALP-induced polarization is small (since the fraction of converted photons is an oscillatory function of distance), it is much less likely that the polarization would be accidentally small for multiple sources at different distances. For this reason, combining sources can help remove islands in parameter space where an ALP signal is undetectable for a specific source.
In Fig. 14 we show the expected region for sensitivity to energy-dependence in the polarization when only a single source is included (for the sources with the highest photon flux), and also for the combined results with that one source removed.
We see that the shape of the constraint region is in both cases generally similar to that of the combined analysis, implying that the signal is not dominated by one source. However, omitting J2021+3651 modestly increases the minimum $g_{a\gamma\gamma}$ that can be probed by the pulsar ensemble, and decreases the mass reach for the (perhaps unphysical) case of a circularly polarized source. The region constrained by W49B alone is markedly narrower in mass than the combined limit. Either omitting J2021+3651 from the pulsar ensemble or including only that source leads to more noticeable “islands” of enhanced detection probability surrounded by regions where a detectable signal is not expected.
Figure 14: Sensitivity to energy-dependence of polarization with 10 years of AMS-100 data, as in the right panel of Fig. 10. Left panel: contours of expected confidence level $95\%$ ($p=0.05$), for several individual Galactic sources. Right panel: contours of expected confidence level $95\%$ ($p=0.05$), for the combined analysis when the specified sources are omitted.

### v2

@@ V.2 Galactic sources
Figure 10: Left panel: region of ALP parameter space where the expected $p$-value satisfies $p<0.05$, for detection of ALP-induced polarization (assuming initially unpolarized sources), using ten years of data from AMS-100P (red) and AMS-100 (blue). The sources employed are four Galactic supernova remnants (see text for details). Right panel: region of ALP parameter space where the expected $p$-value satisfies $p<0.05$, for detection of ALP-induced energy-varying polarization, using ten years of data from AMS-100. Different lines correspond to different source populations, plus (in the case of pulsars) different assumptions for the initial polarization. The two cases with $100\%$ initial linear polarization correspond respectively to the polarization at the source being aligned with the transverse Galactic magnetic field (green dotted line) and the angle between the two being $\approx\tan^{-1}3$ (purple solid line), which maximizes the initial oscillation amplitude.
We summarize the regions of ALP parameter space where we expect to be able to exclude the relevant null hypotheses, using observations of Galactic sources, in Fig. 10. We show the $95\%$ confidence sensitivity to polarization for AMS-100 and AMS-100P in the left panel, and in the right panel show the sensitivity to energy-dependence of polarization for a range of Galactic sources and assumptions about the initial polarization, for AMS-100. In particular, we find that the contours for pulsars and supernova remnants are rather similar, although unpolarized (or circularly polarized) sources allow for probing slightly lower $g_{a\gamma\gamma}$.
@@
Figure 11: Left panel: Region with expected $p$-value satisfying $p\leq 0.05$ for detection of non-vanishing linear polarization, compared to other existing constraints. Right panel: Region with expected $p$-value satisfying $p\leq 0.05$ for detection of energy dependence in linear polarization, compared to other existing constraints. Again, the possible signal from Fermi-LAT [18] is shaded by light blue. The other existing constraints are plotted using AxionLimitPlotter [59].
@@
Similarly, a null result in this search (even with AMS-100) would not nominally constrain new ALP parameter space due to the overlap with existing bounds, but could help test models which evade these bounds due to specific model-dependent features. The uncertainty in the $B$-field modeling is less severe in this case than for the NGC1275 study, but would need to be accounted for in a detailed study of the implications of either a detection or a null result.
@@ V.4 The possibility of a Galactic supernova
The field of view for AMS-02 conversion photons can be estimated as 0.5-0.7 sr (depending mildly on energy), from the results for acceptance and effective area provided in Ref. [27], so the chance that a given Galactic supernova occurs within the field of view is only $\mathcal{O}(5\%)$ (we caution that this is only an order-of-magnitude estimate that does not account for AMS-02’s non-uniform exposure). Nonetheless, we will here take the optimistic approach of assuming a Galactic supernova at a distance of 10 kpc, located such that its photons reach AMS-02 with near-perpendicular incidence. To estimate the counts from such an event, we employ the example axion spectrum arising from pions (which is the dominant contribution), integrated for 10s after the supernova, calculated by Ref. [60] for $g_{a\gamma\gamma}=10^{-12}\text{GeV}^{-1}$ and E=0.11E=0.1-1 GeV. We weight this spectrum by the effective area for AMS-02 pair-conversion photons given in Ref. [27], for E=0.21E=0.2-1 GeV (the effective area falls rapidly to zero below 0.2 GeV). We find that with the “KSVZ” parameters of Ref. [60], we would expect roughly $9.5\times 10^{10}(g_{a\gamma\gamma}/10^{-12}\text{GeV})^{2}$ axions incident on the detector; the measurable photon counts will be this number multiplied by the axion-photon conversion probability. For the “ALP” parameters of that reference (differing from the“KSVZ” case in the couplings relevant for axion production in the supernova), the corresponding number is roughly $26000(g_{a\gamma\gamma}/10^{-12}\text{GeV})^{2}$ axions.
@@
With the red supergiant conversion probabilities, we see that in order to distinguish a fully polarized signal from an unpolarized signal at 95% confidence (i.e. requiring MDP <1<1, so at least $\sim 600$ counts), for AMS-02 we would need $g_{a\gamma\gamma}\gtrsim 7\times 10^{-12}$ GeV-1, for the ALP benchmark and for masses $m_{a}\lesssim 2\times 10^{-11}$ eV (where the conversion in galactic fields dominates). For larger masses, up to $m_{a}\lesssim 5\times 10^{-5}$ eV, the corresponding sensitivity would be $g_{a\gamma\gamma}\gtrsim 3\times 10^{-11}$ GeV-1. For the KSVZ benchmark, the corresponding numbers are $g_{a\gamma\gamma}\gtrsim 2\times 10^{-13}$ GeV-1 ($m_{a}\lesssim 2\times 10^{-11}$ eV) and $g_{a\gamma\gamma}\gtrsim 7\times 10^{-13}$ GeV-1 ($m_{a}\lesssim 5\times 10^{-5}$ eV).
AMS-100 is expected to have a field of view close to $4\pi$, thus greatly improving the odds of seeing the supernova in the first place. If we approximate its effective area as the effective acceptance divided by $4\pi$, we would obtain a peak effective area of $30/4\pi\approx 2.4$ m2, compared to $180$ cm2 for AMS-02. This corresponds to an increase in effective area by a factor of around 130.
Since the axion luminosity and the conversion probability each scale independently as $g_{a\gamma\gamma}^{2}$, an AMS-100-like experiment would improve the sensitivity to $g_{a\gamma\gamma}$ by a factor of $130^{1/4}\approx 3.4$. Similarly, if the stellar conversion probability we use above is underestimated by roughly an order of magnitude (as might be the case for a blue supergiant progenitor with kG surface magnetic field), that would improve the sensitivity to $g_{a\gamma\gamma}$ by a factor of $10^{1/4}\approx 1.8$ in the high-ALP-mass region where conversion on stellar magnetic fields dominates.
@@ VI Summary and Outlook
We have analyzed the detectability of gamma-ray polarization, both astrophysical and induced by new physics such as ALP-photon mixing, in the AMS-02 and (proposed) AMS-100 detectors. Compared with existing X-ray polarization measurements, gamma-ray polarization would both provide a new window on astrophysical production mechanisms for gamma-rays in bright sources, and could offer sensitivity to regions of ALP parameter space where the critical energy for oscillations is high and consequently the signal is suppressed at low photon energies.
We have shown that with AMS-02, we only expect sufficient statistics to detect polarization for the brightest Galactic gamma-ray sources, such as the Vela and Geminga pulsars. These sources are relatively close to Earth and we do not expect a significant ALP-induced polarization signal, but an indication of any gamma-ray polarization in these sources could shed light on their (astrophysical) gamma-ray production mechanisms and serve as a proof of principle for future searches.
Detection of ALP-induced polarization in gamma rays would thus have to wait for a future mission, but we have shown that AMS-100, or a proposed pathfinder with smaller effective area by a factor of 5 (denoted AMS-100P), would each be able to probe currently unconstrained regions of ALP parameter space using observations of the bright AGN NGC1275. Specifically, AMS-100 would extend the $g_{a\gamma\gamma}$ reach down to $\sim 7\times 10^{-13}$ GeV-1 for masses up to $\sim 2\times 10^{-9}$ eV, with the median $B$-field configuration, and would retain much of this sensitivity even for unfavorable $B$-field configurations (at the $5$th percentile of significance, among $B$-field models we tested).
@@
Galactic sources are less favorable for probing small $g_{a\gamma\gamma}$ due to the smaller propagation distance through substantial $B$-fields (compared to NGC1275, which is located in the core of the Perseus cluster). We have shown that the region where AMS-100 has sensitivity is nominally excluded by existing bounds from CAST, SHAFT, and ABRA 10cm. This is broadly consistent with previous studies of energy-dependent modulation of the intensity signal in gamma-rays from pulsars, which claimed a nominal detection of a signal, but in ALP parameter space that is in tension with existing bounds [18]. Some models have been proposed to alleviate or remove these tensions, and in such models, the AMS-100 sensitivity region from our work overlaps with the signal region found in the literature for intensity fluctuations, and could thus serve as an independent check for at least part of the parameter space. Similarly, gamma-ray observations of a Galactic supernova could offer sensitivity to new regions of ALP parameter space [60], and we have estimated the sensitivity of AMS-02 to the polarization signal from such a source (assuming the roughly $5\%$ chance that the supernova occurs within the AMS-02 field of view). AMS-100 would be an excellent supernova detector due to its nearly $4\pi$ field of view, and we have also estimated the sensitivity in this case. Polarization is unlikely to be the first detection channel for ALPs in such an event, as a polarization measurement requires a substantial photon flux (in AMS-like detectors, at least 600 photons are required for a fully polarized signal to be above the MDP with $p=0.05$), but could serve as a powerful cross-check on the ALP origin of such a signal.
@@ Acknowledgements
We thank the anonymous referee for helping us uncover an error in the AMS-02 acceptance modeling in an earlier version of this work. We thank Francesca Calore and Ben Safdi for helpful conversations, and Carlos Maña Barrera for his work on measuring polarization in AMS-02. YJ is grateful for the support of DOE including resources from the National Energy Research Scientific Computing Center under Contract No. DE-AC02- 05CH11231. XZ is supported by DOE grant number DE-SC0024112. TRS was supported in part by a Guggenheim Fellowship; the Edward, Frances, and Shirley B. Daniels Fellowship of the Harvard Radcliffe Institute; the Bershadsky Distinguished Fellowship of the Harvard Physics Department; and the Simons Foundation (Grant Number 929255, T.R.S). TRS thanks the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Mainz Institute for Theoretical Physics for their hospitality during the completion of this work; this research was supported in part by grant no. NSF PHY-2309135 to KITP, and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This work was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567.
@@ Appendix A Note on choice of energy binning
We demonstrate this point in Fig. 12, which shows the sensitivity regions for the energy-dependence analysis for NGC1275 for the cases of 10 bins and 4 bins. We see that when the median-$p$-value $B$-field configuration is chosen for each parameter point (underlying contours), the binning makes very little difference. But when we instead choose an “unfavorable” $B$-field configuration (ranked 95th out of 100 in terms of $p$-value at that point), using only 4 bins removes an $\mathcal{O}(1)$ fraction of the region where the expected $p$-value is below 0.05, in particular erasing sensitivity at lower $m_{a}$, since the “holes” in sensitivity are exactly what dominates the ““unfavorable” case; thus a finer binning improves sensitivity quite considerably. In a real analysis, where there is only one dataset (rather than a large ensemble of simulations), an even finer binning (limited by the AMS-02 energy resolution) might be appropriate.
@@
Figure 12: Sensitivity (1 - expected $p$-value) to energy-dependence of polarization with ten years of AMS-100 data from NGC1275. Contours show the results for the median $B$-field configuration (ranked by expected $p$-value); the cross-hatched region corresponds to $p\leq 0.05$ for an unfavorable $B$-field model (95th/100 in significance). Left panel: results of analysis with 4 log-spaced energy bins. Right panel: results of analysis with 10 log-spaced energy bins, as in Fig. 8.
Figure 13: Sensitivity to energy-dependence of polarization with 10 years of AMS-100 data from NGC1275, as in the bottom right panel of Fig. 8. Left panel: contours of 11- expected $p$-value, for the median $B$-field model. The green line corresponds to the expected $p=0.05$ contour. The cross-hatched region corresponds to expected $p\leq 0.05$ for the 95th best $B$-field scenario (out of 100). Right panel: results of identical analysis accounting for energy resolution effects.
However, as shown in  [27], the energy resolution of AMS-02 is non-negligible: around $15\%$ at 200MeV and grows with energy to $28\%$ at 10 GeV. One might worry that the non-negligible energy resolution will cause a substantial number of photons to be detected in neighboring bins that do not reflect their true energy, thus flattening out the energy dependent polarization pattern, especially for finer binnings. We test this effect explicitly for NGC1275 with 10 bins by assuming an unfavorable $28\%$ resolution for all energy bins. In each energy bin, we randomly draw a thousand samples according to the emission spectrum and, for each sample, we assume the detected photon will follow a Gaussian distribution with a mean value of its true energy and a standard deviation of $28\%$ of the mean value to determine the observed energy. This allows us to model what percentage of photons are transferred to any other bins, as a function of energy. In each energy bin, we average over the polarization as before but now taking into account the transferred photons as well. From there we redo the rest of the analysis as before with the same set of random $B$-field configurations. The result compared to the one neglecting energy-resolution effects is shown in Fig. 13. The plot looks almost identical for $p\lesssim 0.5$, justifying our decision to neglect energy-resolution effects for our default binning (the case with 4 wider bins, rather than 10, will be even less sensitive to the energy resolution), although there is some modification to the contours for higher $p$ (but this in any case corresponds to scenarios where the ALP-induced signal cannot be distinguished from constant polarization).
Appendix B Impact of astrophysical backgrounds
As discussed in the main text, astrophysical diffuse gamma-ray backgrounds may confuse the signal from the sources. In this appendix, we test the effects of backgrounds by adding them to our simulated datasets for AMS-100, assuming that the ALP-induced polarization effect on the backgrounds can be neglected. For each source, we compute the number of expected background counts per bin in energy and azimuthal angle, as described in Sec. IV.1, and then perform a Poisson draw to determine the actual number of background events in each bin in the mock data. We then compute the expected $p$-values (testing for a polarization signal, and for energy-dependent polarization) from these realizations as previously.
@@ Appendix B Impact of astrophysical backgrounds
Figure 14: Contours of $p=0.05$ for testing the presence of polarization (upper panels) and the energy-dependence of polarization (lower panels) with 10 years of AMS-100 data from Galactic sources. Left panels show the results with no background, as in the left and right panels of Fig. 10. Right panels show the results of the identical analysis with background included (in the upper right panel, results are for AMS-100 only).
Fig. 14 shows the impact of this choice on the analyses searching for ALP-induced polarization from Galactic sources. We observe that inclusion of backgrounds noticeably degrades the sensitivity region, largely by removing “islands” of lower sensitivity rather than by changing the minimum testable $g_{a\gamma\gamma}$. Fig. 15 performs the analogous study for NGC1275, where we see there is very little impact on the $p\leq 0.05$ region (although there is some impact on the shape of the contours for higher $p$ in the energy-dependent analysis). This makes sense as the diffuse background is relatively much larger for distant (and therefore fainter) Galactic sources that lie in the Galactic plane, compared to NGC1275, which is both bright and located away from the Galactic plane.
Figure 15: Sensitivity to polarization (upper panels) or energy-dependence of polarization (lower panels) with 10 years of AMS-100 data from NGC1275. Left panels: contours of 11- expected $p$-value, for the median $B$-field model, with no background, as in the right panels of Fig. 8. The green line corresponds to the expected $p=0.05$ contour. The cross-hatched region corresponds to expected $p\leq 0.05$ for the 95th best $B$-field scenario (out of 100). Right panels: results of identical analysis accounting for backgrounds.